# Distributionally Robust Control Synthesis for Stochastic Systems with Safety and Reach-Avoid Specifications

### v1

@@ 2 Preliminary
Notations: We denote by $\mathbb{R}$, $\mathbb{R}_{\geq 0}$ the set of all real numbers and non-negative real numbers, respectively. Given a Borel space $A$, $\mathcal{B}(A)$ and $\mathcal{P}(A)$ represent its Borel $\sigma$-algebra and the set of Borel probability measures on $A$, respectively. Unless otherwise stated, all sets of probability measures are endowed with the weak topology [8, Sec. 7.4.2]. Let $Y$ and $Z$ be separable and metrizable spaces. A stochastic kernel $q(\mathrm{d}z|y)$ on $Z$ given $y$ is a collection of probability measures in $\mathcal{P}(Z)$ parameterized by $y\in Y$. Given $B\in\mathcal{B}(Z)$, we denote by $q(B|y)$ the probability on $B$ under measure $q(\mathrm{d}z|y)$.
@@ 2.1 Discrete-Time Stochastic Control Systems
such that each $\pi_{t}:X\to U$ is measurable. We denote by $\Pi$ the set of control policies. Similarly, an adversary strategy $\gamma$ is a sequence
@@ 2.3 Wasserstein Ambiguity Sets
where $d$ is a metric on $W$ and $\mathbb{M}^{i}(\kappa)$ denotes the $i$-th marginal of $\kappa$ for $i=1,2$. Intuitively, the Wasserstein distance between two probability distributions represents the minimum cost of transporting or redistributing mass from one to the other via non-uniform perturbation, and the optimization variable $\kappa$ can be interpreted as a transport plan [17].
@@ 3 Problem Formulation
As discussed above, in the case of an unknown disturbance, the nominal distribution typically comes from the empirical data. In the case of an antagonistic adversary, $\nu_{N}$ is a user-defined distribution based on prior knowledge and confidence regarding the adversary.
@@ 4.1 Existence of Optimal Control Policies
First, we show that the integration in Eq. (15) is well-defined when sets $G$ and $S$ are compact.
@@
Assume that sets $G$ and $S$ are compact. Then for each $t=0,1,\dots,T$, function $\mathbf{v}_{t}$ is upper semi-continuous, which means that $\mathbf{H}(x,u,\mu,\mathbf{v}_{t})$ is well-defined. Furthermore, we have $\mathbf{v}_{t}(x)\in[0,1]$ for $x\in X$.
@@
In [14], the authors also consider the reach-avoid control synthesis problem within a stochastic game framework. Specifically, they examine the case where the stochastic transition kernels are influenced by the actions of both players. Compared to [14], our result differs in the following ways. First, [14] assumes that for every Borel subset of the state space, the stochastic kernel is continuous w.r.t. to the adversary variable. This assumption does not hold in our setting, as it is inconsistent with convergence of probability measures in the weak topology [8, Sec. 7.4.2]. As a result, while the optimal adversary policy always exists in [14], in our work, as stated in Proposition 2-(2), only a sub-optimal policy always exists for the adversary.
@@ 4.2 Computation Considerations for Optimal Policies
In general, solving the Bellman equation (14) to evaluate $\mathbf{v}_{t}(x)$ with the Wasserstein ambiguity set (11) is challenging, as it involves infinite-dimensional min-max optimization problems. To address this difficulty, we first reformulate the optimization problem in the dynamic programming using duality.
@@
$\int_{W}\inf_{w\in W}\left[\mathbf{v}_{t+1}(f(x,u,w))+\lambda d^{p}(w,\hat{w})\right]\nu_{N}(\mathrm{d}\hat{w})\bigg{\}}$
@@
$v_{D}$ $=\sup_{\lambda\geq 0}\left[-\lambda\theta^{p}+\int_{W}\inf_{w\in W}\left[V(w)+\lambda d^{p}(w,\hat{w})\right]\nu_{N}(\mathrm{d}\hat{w})\right].$
@@
$\int_{W}\inf_{w\in W}\left[\mathbf{v}_{t+1}(f(x,u,w))+\lambda d^{p}(w,\hat{w})\right]\nu_{N}(\mathrm{d}\hat{w})\bigg{\}}.$
@@
[…]
Intuitively, when the nominal distribution has finite support, Eq.(30) replaces the integral in Eq.(27) with a probabilistic weighted sum, where each variable $l_{i}$ represents the infimum value when $\hat{w}=\hat{w}_{i}$. The optimization problem in Eq. (30) is a semi-infinite optimization program for which various convergent methods exist, such as discretization methods, primal-dual methods, and constraint sampling methods [18, 21, 24]. In our later experiments, we employ the discretization algorithm from [32], which adaptively generates grids over $W$. This approach ensures convergence to a local optimal solution of the semi-infinite program.
@@ 4.3 Case of Safety Specifications
$\int_{W}\inf_{w\in W}\left[\hat{\mathbf{v}}_{t+1}(f(x,u,w))+\lambda d^{p}(w,\hat{w})\right]\nu_{N}(\mathrm{d}\hat{w})\bigg{\}}$
@@ 5.2 Searching DR-CBC via SOS Programs
    Each set $\star\in\{X_{0},X,S,X\setminus S,S\times U\times W\}$ is defined as the super-level set of polynomial $\mathbf{s}_{\star}$, i.e., =$\star=\{x\in\mathbb{R}^{n}\mid\mathbf{s}_{\star}(x)\geq 0\}$. For simplicity, we denote $\mathbf{s}_{S\times U\times W}(x)$ as $\mathbf{s}_{\zeta}(x)$;
@@
$-\theta^{p}\lambda(x)+\sum_{i=1}^{M}p_{i}l_{i}(x)-\frac{\tilde{\mathbf{v}}(x)}{\eta}\geq\beta,\forall x\in S.$ (56)
@@
In Theorem 4, we determine $\tilde{\mathbf{v}}$ using SOS programs under Assumption 3. However, for more general nonlinear systems, the program (46)-(52) cannot be directly converted into an SOS program. Several potential approaches exist to address such cases. One possibility is to utilize satisfiability modulo theories (SMT) solvers within a counterexample-guided inductive synthesis framework to find $\tilde{\mathbf{v}}$; see [5, 12, 28]. Another alternative is to parameterize $\tilde{\mathbf{v}}$ using a neural network, enabling the synthesis of neural certificates based on constraints (46)-(52); see, for example, [15].
@@ 6 Case Studies
In this section, we illustrate our methods with two case studies. The first case study addresses the reach-avoid specification, employing the duality-based approach. The associated optimization problem is solved using Gurobi [20]. The second case study focuses on the safety specification, utilizing the DR-CBC approach. The SOS program involved is solved with SOSTOOLS [30].
@@ 6.1 Room Temperature Control
Table 1: Statistic results for $5$ groups of parameters
[b] Parameter group $1$ $2$ $3$ $4$ $5$ ave. $\mathsf{RA}$ prob. of us $0.96$ $0.96$ $0.96$ $0.96$ $0.96$ success rate of us $1$ $1$ $1$ $1$ $1$ ave. $\mathsf{RA}$ prob.  of [38] $0.59$ $0.78$ $0.87$ $0.88$ $0.95$ success rate of [38] $0.5$ $0.6$ $0.7$ $0.6$ $1$
Table 2: Detailed experimental results of the $2$-nd group of parameters $N=5$ and $\theta=0.05$ using [38]
[b] Round $1$ $2$ $3$ $4$ $5$ $6$ $7$ $8$ $9$ $10$ $\mathsf{RA}$ probability $0.94$ $0.96$ $0.82$ $0.56$ $0.96$ $0.96$ $0.96$ $0.35$ $0.96$ $0.35$ Distribution mean $0.0019$ $-0.022$ $0.0018$ $0.043$ $-0.025$ $-0.0036$ $-0.03$ $0.047$ $-0.018$ $0.037$
@@
where $T_{e}=15$, $T_{h}=50$, $\alpha_{e}=8\times 10^{-3}$, $\alpha_{h}=3.6\times 10^{-3}$ and $\tau_{s}=5$. The initial state set, control input set and disturbance set are defined as $X_{0}=[23.6,23.8]$, $U=\{0,1\}$ and $W=[-0.12,0.12]$, respectively. We assume that at each time step, $w_{t}$ follows a uniform distribution over $W$, which is unknown a priori. The control objective is that, within a time horizon of $T=12$, the room temperature should reach the target region $G=[24.4,24.6]$ while staying within the safe region $S=[23,26]$ with a probability threshold $\alpha=0.9$.
@@
For each group of parameters, we repeat the experiment 10 times. In each iteration, we solve the reach-avoid synthesis problem using both our distributionally robust approach (with respect to the Wasserstein ball defined by $\nu_{N}$ and $\theta$) and the stochastic optimal control approach from [38], which relies on the empirical distribution $\nu_{N}$. When applying each strategy, to reduce energy consumption, the controller always opts not to heat (i.e., $u=0$) whenever $0\in\mathcal{U}_{t}(x)$ in Eq. (31). To evaluate the performances of the synthesized policies, we discretize the initial state set $X_{0}=[23.6,23.8]$ with a resolution of $0.01$. For each initial state, we explicitly compute the reach-avoid probabilities under the actual distribution, which is unknown a priori. We then record the smallest reach-avoid probability across all $20$ discretized initial states. The experimental results are presented in Table 1.
@@
Specifically, the first and third rows of Table 1 show the average (across 10 repeated experiments) of the smallest reach-avoid probabilities for each parameter group using our method and the method in [38], respectively. The second and fourth rows record the success rate of each method. A synthesized policy is considered successful if the smallest reach-avoid probability exceeds the given threshold $\alpha=0.9$. Our method ensures the desired satisfaction probability, achieving a success rate of 100%. Note that the reason all parameter groups yield the same satisfaction probability with our method is that this is already the optimal satisfaction probability achievable for this system under the actual uniform distribution. In contrast, the stochastic control framework based solely on the empirical distribution yields significantly lower average satisfaction probabilities and success rates, particularly when the sample size is insufficient.
@@
To illustrate this more explicitly, we present the details of the ten repeated experiments for the second group of parameters ($N=5$ and $\theta=0.05$) using the stochastic control approach without considering robustness. The results show that even when the sampled mean deviates only slightly from the actual distribution, the resulting controller can exhibit poor satisfaction probabilities. In contrast, our distributionally robust framework achieves a 100% success rate by incorporating a small robustness radius $\theta=0.05$. It is worth noting that this 100% success rate is not guaranteed by our formal analysis, as our theoretical guarantees rely on the correctness of the ambiguity set. Nevertheless, the experimental results demonstrate that accounting for the worst-case scenarios during offline synthesis leads to consistently strong empirical performance in online applications.
6.2 Two-Tank Systems using DR-CBC
@@
To illustrate the method of DR-CBC, we consider the two-tank system control problem adopted from  [31] for safety specifications. The discretized system dynamic with $\tau_{s}=1$ is
$h_{1}(t+1)$ $=(1-\tau_{s}\frac{b_{1}}{a_{1}})h_{1}(t)+\tau_{s}\frac{u(t)+w(t)}{a_{1}}$
$h_{2}(t+1)$ $=\tau_{s}\frac{b_{1}}{a_{2}}h_{1}(t)+(1-\tau_{s}\frac{b_{2}}{a_{2}})h_{2}(t),$
@@
where $[h_{1}(t);h_{2}(t)]\in\mathbb{R}^{2}$ and $w(t),u(t)\in\mathbb{R}$ are state, disturbance and control input at time instant $t$, respectively. The initial state set, control input set, and disturbance set are $X_{0}=[4.5,5.5]\times[4.5,5.5]$, $U=[1,4]$, and $W=[-0.4,0.4]$, respectively. The tank parameters are $a_{i}=10.2$ and $b_{i}=0.77$ for each $i\in\{1,2\}$. Additionally, the true but unknown distribution of $w(t)$ is a truncated Gaussian distribution over $W$ with a standard deviation of 0.4. The overall control objective is to ensure that, within a time horizon of $T=40$, the heights of the tanks remain within the region $S=[1,9]\times[1,9]$ with a probability exceeding the threshold $\alpha=0.9$.
@@
We take $N=5$ empirical samples from the underlying unknown distribution to define the nominal distribution $\nu_{N}$. The order of Wasserstein metric is $p=2$ and radius of the ambiguity set is $\theta=0.05$. In the experiment, the collected sample set is $[-0.2,0.24,0.02,-0.16,0.31]$. We use the SOS program (46)-(52) to solve the control synthesis for safety specification. The degree of polynomial is selected to be $4$ and we choose $\eta=1$, $\beta=-0.0015$ and $\delta=0.96$, which ensure that $\delta+T\beta=0.9$.
The DR-CBC $\tilde{\mathbf{v}}(h_{1},h_{2})$ and its associated control function $\mathbf{u}(h_{1},h_{2})$ computed are
$\tilde{\mathbf{v}}(h_{1},h_{2})=$ $-0.005895h_{1}^{4}+0.005458h_{1}^{3}h_{2}+0.06884h_{1}^{3}$
$-0.004294h_{1}^{2}h_{2}^{2}-0.03513h_{1}^{2}h_{2}-0.3582h_{1}^{2}$
$+0.00685h_{1}h_{2}^{3}-0.05847h_{1}h_{2}^{2}+0.4335h_{1}h_{2}$
$+0.5924h_{1}-0.00559h_{2}^{4}+0.06404h_{2}^{3}$
$-0.2885h_{2}^{2}+0.3351h_{2}-1.051,$
$\mathbf{u}(h_{1},h_{2})=$ $-2.593e^{-4}h_{1}^{4}+6.776e^{-5}h_{1}^{3}h_{2}+0.004289h_{1}^{3}$
$-2.784e^{-5}h_{1}^{2}h_{2}^{2}-6.626e^{-4}h_{1}^{2}h_{2}-0.0275h_{1}^{2}$
$+2.73e^{-4}h_{1}h_{2}^{3}-0.00347h_{1}h_{2}^{2}+0.0191h_{1}h_{2}$
$+0.05536h_{1}-0.0001667h_{2}^{4}+0.00165h_{2}^{3}$
$-0.00299h_{2}^{2}-0.01954h_{2}+3.951.$
@@
Then we simulate the system using the synthesized control policy $\pi_{\mathbf{u}}$ for 1000 trials, under the unknown truncated Gaussian distribution. In all 1000 simulations, the system remains within the safe region. Although our theoretical results guarantee a satisfaction probability of 0.9, this is established for all possible distributions within the ambiguity set. In practice, our simulations achieve a 100% success rate because the applied robust policy effectively accommodates the actual underlying distribution, which is not necessarily the worst-case scenario.

### v2

@@ 2 Preliminary
Notations: We denote by $\mathbb{R}$, $\mathbb{R}_{\geq 0}$ the set of all real numbers and non-negative real numbers, respectively. Given a Borel space $A$, $\mathcal{B}(A)$ and $\mathcal{P}(A)$ represent its Borel $\sigma$-algebra and the set of Borel probability measures on $A$, respectively. Unless otherwise stated, all sets of probability measures are endowed with the weak topology [9, Sec. 7.4.2]. Let $Y$ and $Z$ be separable and metrizable spaces. A stochastic kernel $q(\mathrm{d}z|y)$ on $Z$ given $y$ is a collection of probability measures in $\mathcal{P}(Z)$ parameterized by $y\in Y$. Given $B\in\mathcal{B}(Z)$, we denote by $q(B|y)$ the probability on $B$ under measure $q(\mathrm{d}z|y)$. For simplicity we write Borel measurable directly as measurable.
@@ 2.1 Discrete-Time Stochastic Control Systems
such that each $\pi_{t}:X\to U$ is measurable. Although the policy should generally be universally measurable to preserve measurability in dynamic programming [9, Page 156], a Borel measurable policy is sufficient in our work since the value functions considered are always upper semi-continuous. We denote by $\Pi$ the set of control policies. Similarly, an adversary strategy $\gamma$ is a sequence
@@ 2.3 Wasserstein Ambiguity Sets
where $d$ is a metric on $W$ and $\mathbb{M}^{i}(\kappa)$ denotes the $i$-th marginal of $\kappa$ for $i=1,2$. Intuitively, the Wasserstein distance between two probability distributions represents the minimum cost of transporting mass from one to the other via non-uniform perturbation, and the optimization variable $\kappa$ can be interpreted as a transport plan [19]. The Wasserstein distance offers several advantages, such as leveraging more information from the nominal distribution than the moment-based approach, and encompassing a wider range of relevant distributions compared to many $\phi$-divergences, such as Kullback-Leibler divergence. For further details on the advantages of the Wasserstein ambiguity set, readers are referred to [19].
@@ 3 Problem Formulation
One may regard that the adversary player should use a stationary policy when it is actually an unknown disturbance. However, using a time-invariant adversary setting will lead to an intractable program since the original decision problem involves multi-step dynamic system evolution with decision variables at each step. This setting intends to compress the complex multi-stage decision problem into a high-dimensional single-stage problem. The high-dimensional issue must be addressed to adopt a less conservative setting.
@@
As discussed above, in the case of an unknown disturbance, the nominal distribution typically comes from the empirical data. In the case of an antagonistic adversary, $\nu_{N}$ is a user-defined distribution based on prior knowledge and confidence regarding the adversary. Note that even though the nominal distribution is finite-support, the Wasserstein ambiguity set still contain continuous distribution since the Wasserstein distance is possible to be finite between discrete and continuous distributions.
@@ 4.1 Existence of Optimal Control Policies
Intuitively, $\mathbf{v}_{0}(x)$ characterizes the reach-avoid probability starting from state $x\in X$ under optimal control policy and adversary strategy. First, we show that the integration in Eq. (15) is well-defined when sets $G$ and $S$ are compact.
@@
Assume that sets $G$ and $S$ are closed. Then for each $t=0,1,\dots,T$, function $\mathbf{v}_{t}$ is upper semi-continuous, which means that $\mathbf{H}(x,u,\mu,\mathbf{v}_{t})$ is well-defined. Furthermore, we have $\mathbf{v}_{t}(x)\in[0,1]$ for $x\in X$.
@@
In [16], the authors also consider the reach-avoid control synthesis problem within a stochastic game framework. Although [16] addresses a more general stochastic hybrid system than our continuous system, the dynamic game framework in [16] is inapplicable to our work. Specifically, [16] assumes that for every Borel subset of the state space, the stochastic kernel is continuous with respect to the adversary variable. This assumption does not hold in our setting, as it conflicts with the convergence of probability measures in the weak topology [9, Sec. 7.4.2]. Moreover, due to the differences in game frameworks, while the optimal adversary policy always exists in [16], our work (as stated in Proposition 2-(2)) guarantees only the existence of a sub-optimal policy for the adversary.
@@ 4.2 Computation Considerations for Optimal Policies
In general, solving the Bellman equation (14) to evaluate $\mathbf{v}_{t}(x)$ with the Wasserstein ambiguity set (11) is challenging, as it involves infinite-dimensional min-max optimization problems. To address this difficulty, we first reformulate the optimization problem in the dynamic programming for each stage of value function using duality result in [19].
@@
[…]
@@
[…]
@@
[…]
@@
$\text{s.t.}\quad l_{i}\leq\mathbf{v}_{t+1}(f(x,u,w))+\lambda d^{p}(w,\hat{w}_{i}),\forall w\in W.$
Intuitively, when the nominal distribution has finite support, Eq.(30) replaces the integral in Eq.(27) with a probabilistic weighted sum, where each variable $l_{i}$ represents the infimum value when $\hat{w}=\hat{w}_{i}$. Note that the variables $\lambda$ and $u$ in Eq. (30) may have multiple optimal solutions. Moreover, every optimal solution $u$ corresponds to the optimal control policy at time instant $t$ and state $x$, i.e., the policy $\pi_{t}(x)$. The optimization problem in Eq. (30) is a semi-infinite optimization program for which various convergent methods exist, such as discretization methods, primal-dual methods, and constraint sampling methods [20, 23, 26]. In our later experiments, we employ the discretization algorithm from [34], which adaptively generates grids over $W$. This approach ensures convergence to a local optimal solution of the semi-infinite program. Note that the complexity of computing $\mathbf{v}_{t}$ is exponential in the dimension of $X$, as solving via dynamic programming requires discretizing $X$. Additionally, in the more complex distributionally robust setting, when solving Eq. (30) for a specific $x\in X$, the number of constraints typically grows exponentially with the dimension of the disturbance space $W$. This complexity is difficult to avoid, as the optimization involves general non-convex and potentially discontinuous functions $\mathbf{v}_{t}$. To alleviate this exponential complexity, in the next section, we will demonstrate how to compute an approximate solution to the safety control synthesis problem.
@@
Remark 3.
Here we discuss how $\theta$ the satisfaction probability, which is equivalent to the objective value of Eq. (30). When $\theta=0$, the Wasserstein ambiguity set only contains the distribution $\sum_{i=1}^{M}p_{i}\delta_{\hat{w}_{i}}$. In this case, the optimal value of Eq. (30) $\sum_{i=1}^{M}p_{i}\mathbf{v}_{t+1}(f(x,u,w_{i})), si$nce $\lambda$ should be set to infinity. When $\theta=\infty$, the Wasserstein ambiguity set contains all probability measures on $W$, and the optimal value of Eq. (30) is $\inf_{w\in W}\{\mathbf{v}_{t+1}(f(x,u,w))\}$, since $\lambda$ should be set to zero in this case. Therefore, as $\theta$ increases, the Wasserstein ambiguity set includes more distributions, and the computed reach-avoid probability becomes smaller. We refer readers to [19] for more details on how to choose $\theta$ such that the true distribution is contained within the ambiguity set $\mathbb{D}$ with a given confidence, while ensuring that the distributionally robust optimization is not overly conservative.
@@ 4.3 Case of Safety Specifications
[…]
@@ 5.2 Searching DR-CBC via SOS Programs
    Each set $\star\in\{X_{0},X,S,X\setminus S,S\times U\times W\}$ is defined as the super-level set of polynomial $\mathbf{s}_{\star}$, i.e., $\star=\{x\in\mathbb{R}^{n}\mid\mathbf{s}_{\star}(x)\geq 0\}$. For simplicity, we denote $\mathbf{s}_{S\times U\times W}(x)$ as $\mathbf{s}_{\zeta}(x)$;
@@
$\theta^{p}\lambda(x)+\sum_{i=1}^{M}p_{i}l_{i}(x)-\frac{\tilde{\mathbf{v}}(x)}{\eta}\geq\beta,\forall x\in S.$ (56)
@@
In Theorem 4, we determine $\tilde{\mathbf{v}}$ using SOS programs under Assumption 3. However, for more general nonlinear systems, the program (46)-(52) cannot be directly converted into an SOS program. Several potential approaches exist to address such cases. One possibility is to utilize satisfiability modulo theories (SMT) solvers within a counterexample-guided inductive synthesis framework to find $\tilde{\mathbf{v}}$; see [6, 14, 31]. Another alternative is to parameterize $\tilde{\mathbf{v}}$ using a neural network, enabling the synthesis of neural certificates based on constraints (46)-(52); see, for example, [27, 17]. Note that it is non-trivial to search for DR-CBC using SMT solvers or neural networks since Eq. (37) involves the computation of the expectation of system evolution under a distributionally robust setting, and thus we regard it as a future direction.
@@ 6 Case Studies
In this section, we illustrate our methods with three case studies. The first case study addresses the reach-avoid specification, employing the duality-based approach. The associated optimization problem is solved using Gurobi [22]. The second and third case study focuses on the safety specification, utilizing the DR-CBC approach. The SOS program involved is solved with SOSTOOLS [33] and Mosek solver [3].
@@ 6.1 Room Temperature Control
Table 1: Statistic results for 55 groups of parameters
[b] Parameter group 11 22 33 44 55 ave. $\mathsf{RA}$ prob. of us (%) $95.7$ $95.8$ $95.7$ $95.8$ $95.8$ success rate of us (%) $99$ $100$ $99$ $100$ $100$ ave. $\mathsf{RA}$ prob. of [40] (%) $66.2$ $85.5$ $88.8$ $89.5$ $93.0$ success rate of [40] (%) $54$ $66$ $68$ $74$ $86$
Table 2: Detailed experimental results of the 22-nd group of parameters $N=5$ and $\theta=0.05$ using [40]
[b] Round 11 22 33 44 55 66 77 88 99 $10$ $\mathsf{RA}$ probability $0.94$ $0.96$ $0.82$ $0.56$ $0.96$ $0.96$ $0.96$ $0.35$ $0.96$ $0.35$ Distribution mean $0.0019$ 0.022-0.022 $0.0018$ $0.043$ 0.025-0.025 0.0036-0.0036 0.03-0.03 $0.047$ 0.018-0.018 $0.037$
@@
where $T_{e}=15$, $T_{h}=50$, $\alpha_{e}=8\times 10^{-3}$, $\alpha_{h}=3.6\times 10^{-3}$ and $\tau_{s}=5$. The initial state set, control input set and disturbance set are defined as $X_{0}=[23.6,23.8]$, $U=\{0,1\}$ and W=[0.12,0.12]W=[-0.12,0.12], respectively. We assume that at each time step, $w_{t}$ follows a uniform distribution over $W$, which is unknown a priori. The control objective is that, within a time horizon of $T=12$, the room temperature should reach the target region $G=[24.4,24.6]$ while staying within the safe region $S=[23,26]$ with a probability threshold $\alpha=0.9$.
@@
For each group of parameters, we repeat the experiment 100 times. In each iteration, we solve the reach-avoid synthesis problem using both our distributionally robust approach (with respect to the Wasserstein ball defined by $\nu_{N}$ and $\theta$) and the stochastic optimal control approach from [40], which relies on the empirical distribution $\nu_{N}$. When applying each strategy, to reduce energy consumption, the controller always opts not to heat (i.e., $u=0$) whenever $0\in\mathcal{U}_{t}(x)$ in Eq. (31). To evaluate the performances of the synthesized policies, we discretize the initial state set $X_{0}=[23.6,23.8]$ with a resolution of $0.01$. For each initial state, we explicitly compute the reach-avoid probabilities under the actual distribution, which is unknown a priori. We then record the smallest reach-avoid probability across all $20$ discretized initial states. The experimental results are presented in Table 1.
@@
Specifically, the first and third rows of Table 1 show the average (across 100 repeated experiments) of the smallest reach-avoid probabilities for each parameter group using our method and the method in [40], respectively. The second and fourth rows record the success rate of each method. A synthesized policy is considered successful if the smallest reach-avoid probability exceeds the given threshold $\alpha=0.9$. Our method ensures the desired satisfaction probability in almost all trials. Note that the reason all parameter groups yield close satisfaction probability with our method is that every successful trial gets the optimal satisfaction probability achievable for this system under the actual uniform distribution. In contrast, the stochastic control framework based solely on the empirical distribution yields significantly lower average satisfaction probabilities and success rates, particularly when the sample size is insufficient.
@@
To illustrate this more explicitly, we present the details of ten of repeated experiments for the second group of parameters ($N=5$ and $\theta=0.05$) using the stochastic control approach without considering robustness. The results show that even when the sampled mean deviates only slightly from the actual distribution, the resulting controller can exhibit poor satisfaction probabilities. In contrast, our distributionally robust framework achieves a 100% success rate by incorporating a small robustness radius $\theta=0.05$. It is worth noting that this 100% success rate is not guaranteed by our formal analysis, as our theoretical guarantees rely on the correctness of the ambiguity set. Nevertheless, the experimental results demonstrate that accounting for the worst-case scenarios during offline synthesis leads to consistently strong empirical performance in online applications.
6.2 $1$-D System Control using DR-CBC
@@
To illustrate the method of DR-CBC, in this subsection, we consider the one dimension system control problem for safety specifications. The system dynamic is
$x_{t+1}=x_{t}+0.1x_{t}^{2}+u_{t}+w_{t}$
@@
where $x_{t}\in\mathbb{R}$ and $w_{t},u_{t}\in\mathbb{R}$ are state, disturbance and control input at time instant $t$, respectively. The initial state set, control input set, and disturbance set are $X_{0}=[-0.5,0]$, $U=[0,2]$, and W=[4,1]W=[-4,1], respectively. Additionally, the true but unknown distribution of $w(t)$ is a truncated Gaussian distribution over $W$ with a standard deviation of 2. The overall control objective is to ensure that, within a time horizon of $T=40$, the system state is within the region $S=\{x\in\mathbb{R}\mid x\geq-2\}$ with a probability exceeding the threshold $\alpha=0.9$.
@@
We take $N=5$ empirical samples from the underlying unknown distribution to define the nominal distribution $\nu_{N}$. The order of Wasserstein metric is $p=2$, and radius of the ambiguity sets are $\theta_{1}=0.1$ and $\theta_{2}=0.01$. We use the SOS program (46)-(52) to solve the control synthesis for safety specification. The degree of polynomial is selected to be $4$ and we choose $\eta=1$, $\beta=-0.0015$ and $\delta=0.96$, which ensure that $\delta+T\beta=0.9$. The CBCs are computed successfully by the SOS programs under both $\theta_{1}$ and $\theta_{2}$, which can be found in [12].
@@
We simulate the system using $\pi_{\mathbf{u}}^{1}$ and $\pi_{\mathbf{u}}^{2}$ for 10000 trials, under the unknown truncated Gaussian distribution such that $\pi_{\mathbf{u}}^{i}$ is the synthesized control policy under radius $\theta_{i}$. The success rates of safety specification in simulation under $\pi_{\mathbf{u}}^{1}$ and $\pi_{\mathbf{u}}^{2}$ are $99.3\%$ and $91.8\%$, respectively. Although our theoretical results guarantee a satisfaction probability of 0.9, this is established for all possible distributions within the ambiguity set. In practice, our simulations achieve a higher success rate because the applied robust policy effectively accommodates the actual underlying distribution, which is not necessarily the worst-case scenario. Moreover, when the radius is larger, more distributions are included in ambiguity set, and thus the computed control policy is more conservative. Therefore, the success rate under $\pi_{\mathbf{u}}^{1}$ is higher than that under $\pi_{\mathbf{u}}^{2}$.
6.3 $4$-D System Control using DR-CBC
Finally, we consider safety control for $4$-D polynomial system in [43] to show the scalability of proposed DR-CBC and SOS program. The system dynamic is
$x^{1}_{t+1}$ $=x^{1}_{t}+\tau_{s}(-x^{1}_{t}+(x^{2}_{t})^{3}-3x^{3}_{t}x^{4}_{t}+u_{t}+w_{t})$
$x^{2}_{t+1}$ $=x^{2}_{t}+\tau_{s}(-x^{1}_{t}-(x^{2}_{t})^{3})$
$x^{3}_{t+1}$ $=x^{3}_{t}+\tau_{s}(x^{1}_{t}x^{4}_{t}-x^{3}_{t})$
$x^{4}_{t+1}$ $=x^{4}_{t}+\tau_{s}(x^{1}_{t}x^{3}_{t}-(x^{4}_{t})^{3})$
where sample time $\tau_{s}=0.01$, $x_{t}^{1},x_{t}^{2},x_{t}^{3},x_{t}^{4}]\in\mathbb{R}^{4}$ and $w_{t},u_{t}\in\mathbb{R}$ are state, disturbance and control input at time instant $t$, respectively. The initial state set, control input set, and disturbance set are $X_{0}=\{x\in\mathbb{R}^{4}\mid x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{4}^{2}\leq 0.09\}$, U=[1,1]U=[-1,1], and W=[0.8,0.8]W=[-0.8,0.8], respectively. Additionally, the true but unknown distribution of $w(t)$ is a truncated Gaussian distribution over $W$ with a standard deviation of $0.8$. The overall control objective is to ensure that, within a time horizon of $T=100$, the system state remains within the region $S=\{x\in\mathbb{R}^{4}\mid x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{4}^{2}\leq 1\}$ with a probability exceeding the threshold $\alpha=0.92$.
We take $N=5$ empirical samples from the underlying unknown distribution to define the nominal distribution $\nu_{N}$. The order of Wasserstein metric is $p=2$ and radius of the ambiguity set is $\theta=0.1$. We use the SOS program (46)-(52) to solve the control synthesis for safety specification. The degree of polynomial is selected to be $4$ and we choose $\eta=1$, $\beta=-0.0004$ and $\delta=0.96$, which ensure that $\delta+T\beta=0.92$. The computed CBC can be found in [12]. Then we simulate the system using the synthesized control policy $\pi_{\mathbf{u}}$ for 10000 trials, under the unknown truncated Gaussian distribution. The success rate of safety specification in simulation under $\pi_{\mathbf{u}}$ is $98.13\%$, which achieves the required satisfaction probability $0.92$.